OpenStudy (anonymous):

Solve for x: -2|x - 3| = -12

OpenStudy (unklerhaukus):

-2|x - 3| = -12 |x-3| = 6 x= -3 , 6

OpenStudy (anonymous):

It's a multi-choice question... It's not that answer... a) x = -3, x = 9 b) x = 3, x = 9 c) x = -3, x = 4.5 d) No solutions

OpenStudy (unklerhaukus):

ops i cant count the answer is x = -3, 9 that is a)

OpenStudy (anonymous):

The meaning of |x-3| = 6 is the x-values that are exactly 6 units from 3, ie,So you convert the absolute values equation to\[x-3=\pm6\]or \[x=3\pm6\]which gives\[x=3-6=-3\]or\[x=3+6=9\]

OpenStudy (anonymous):

And can you guys help me with these two final questions? :"c

OpenStudy (anonymous):

Solve 2|5x - 2| + 4 = 18 for x. Show your work to receive full credit.

OpenStudy (unklerhaukus):

2|5x - 2| + 4 = 18 |5x - 2| + 2 = 9 |5x - 2| = 7

OpenStudy (anonymous):

and this one! c: Create an absolute value equation that has two solutions. Explain why this equation has two solutions.

OpenStudy (anonymous):

2|5x - 2| + 4 = 18 2|5x - 2| =14 subtract 4 |5x - 2| =7 divide by 2 5x-2=-7 or 5x-2=7 definition of abs value 5x=-5 or 5x=9 add 2 x=-1 or x=9/5 divide by 5 {-1, 9/5} solution set

OpenStudy (unklerhaukus):

...follow Mandolino's methods

OpenStudy (anonymous):

and this one! c: Create an absolute value equation that has two solutions. Explain why this equation has two solutions. any abs value equation set equal to a positive number, eg, |x-5|=7 Has two solutions because the interpretation is the x-values that are exactly 7 units from 5. of course there are two-- one in each direction: